OpenStudy (anonymous):

Find the derivative of f(x) =-11/x at x = 9.

OpenStudy (michele_laino):

please set x= 9 into your formula for f(x)

OpenStudy (michele_laino):

oops.. sorry I have made an error

OpenStudy (michele_laino):

please what is the first derivative of f(x) = -1/x ?

OpenStudy (michele_laino):

oops... of f(x) = -11/x ?

OpenStudy (anonymous):

oh wait I get it

rvc (rvc):

f(x)= -11/x

OpenStudy (anonymous):

it would be 11/81

OpenStudy (anonymous):

becuase 9*9=81

OpenStudy (michele_laino):

please do you know what is the formula of the first derivative of f(x) = -11/x?

rvc (rvc):